OpenStudy (anonymous):

help Find the measure of angle CDE. (Points : 2) 50° 65° 115° 130°

OpenStudy (anonymous):

@jayz657

OpenStudy (nurali):

130

OpenStudy (anonymous):

@jayz657 is tht right?

OpenStudy (anonymous):

@Nurali are u sure? cuz for number 1 u didnt know so im kind scare

OpenStudy (anonymous):

180 - 2x + 15 = 130 - x x = 65 let me look at this a bit longer

OpenStudy (anonymous):

oik

OpenStudy (anonymous):

yea i think its 65

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Nurali u see its 65! -__-